OpenStudy (anonymous):

[(sqrt 3)/2 + 1/2] ^2

OpenStudy (daniellelovee):

are you familiar with fractions addition?

OpenStudy (solomonzelman):

yes, you shall add the fractions. your steps are: 1. Add the fractions inside the brackets. (they have common denominator already) 2. Use the following rule: \(\large\color{slate}{ \displaystyle \left[\frac{{\rm \color{red}{a}}}{{\rm \color{blue}{b}}}\right]^2=\frac{{\rm \color{red}{a}}^2}{{\rm \color{blue}{b}}^2} }\) 3. Simplify the result as much as you can.

OpenStudy (jdoe0001):

\(\bf \left(\cfrac{\sqrt{3}}{2}+\cfrac{1}{2}\right)^2\implies \left(\cfrac{\sqrt{3}}{2}\right)^2+\left(\cfrac{1}{2}\right)^2\implies \cfrac{(\sqrt{3})^2}{(2)^2}+\cfrac{(1)^2}{(2)^2}\) as suggested by SolomonZelman so add them up

OpenStudy (anonymous):

Okay the answer is \[\frac{ 2+\sqrt{3} }{ 2 }\] How do I find that?